OpenStudy (anonymous):

x/9+1/6=2/3 those are fractions

sam (.sam.):

(x)/(9)+(1)/(6)=(2)/(3) x/9=1/2 x=9/2

OpenStudy (anonymous):

uhh y did u not multiply by 9

sam (.sam.):

I multiply both sides by 9

OpenStudy (anonymous):

ok thanks